Dublin University Rifle Club

Version Date : April 24, 2002 Filename : Status :

Dublin University Rifle Club Match Secretary’s Report

Mark Dennehy Match Secretary

April 24, 2002

1 Summary

We attended 19 external competitions this year, down on last year’s 24, but we did achieve higher scores, so we did have an improvement overall. DURC shooters competed in 3 international competitions, and a DURC shooter was the Head of Delegation in the Munich World Cup. Records are not totally complete, but as shown in the table, we have at least 142 entries in various competitions this year, and took away 13 medals and 44 prizes in total. This is quite impressive for one club. We also saw a return to Bisley this year, and hopefully this marks a resumption of the traditional involvement in competitions in the UK and abroad.

2 DURC competitions hosted

Foot and Mouth played havoc with last year’s calendar, and this year’s calendar seems not to have settled yet. As a result, we have not yet held the 2002 50m Open, but have the result of the 2001 50m Open instead.

2.1 DURC 10m Air Rifle Open

The DURC 10m Open maintained the high level of excellence that was set in the 2000-1 academic year. We had 35 competitors and 43 entries. All finalists shot 550 or above. Mark and Steven made it to the finals, Steven taking third place. The finals had three close competitions with tied scores, on first place, third place and fifth place. Nicola Murphy won, despite a misfire costing her nine points. Simon Kennedy donated a large sum towards prizes and as a result we now have the Simon Kennedy Air Rifle Shield for the next few years as

2 DURC competitions hosted

Competition  Entries  Prizes Won
Overall  Class
A B C D
10m 50m 25yd 1 2 3 1 2 3 1 2 3 1 2 3 1 2 3
DURC 10m Open 12  1  1
DURC 50m Open  7  1 1  1 1  1  1 1
EARC Festival 2  3
FSC 50m Open  7  1
Island Games 1  1  2 1
NSRA Postal 2  4
Leinster Indoor  9  1  1  1  1
UCD Open 1 10  1  1  1  1 1
UCD Open 2 6  1
FSC 10m Open 1 10  1  1
WTSC Open 1 4  1  1  1  1
WTSC Open 2 3
Air Nationals 2001 3  1
Air Nationals 2002 24
50m Nationals 2001  9  1  1  1 1 1  1
25yd Nationals 2002  9 1  1  1 1  1 1 1
Comber Open 2001  5
IoM Easter Shoot 2 3 1
Bisley Open 6
85  35  22 3 5 5  1 3 4 1 1 3 4 1 5 3 5
Totals  142  13  4  6  8  13

2 a perpetual prize. Hammereli declined to sponsor us, but did send lots of caps which were given out to all the juniors.

2.2 DURC 50m Open

The DURC 50m Open was sponsored by Shilen Barrels, who gave us a voucher for a custom rifle barrel. This was quite a valuable prize and was given out for first place. Steven kept the prize in the club with a 585, and Ray came in second with 582 to ensure the barrel stayed with us. We also held a 50m rapid-fire competition, which proved quite novel, and which was won by Ray with a 97/100. Steven and Alan Lewis had a shoot-off for second place, which Steven won. The barbecue was a success, but this year instead of too much vegetarian food, we had too much meat. This year we must ensure that we include the other FSC shooters in the barbecue or we’ll have food being wasted again.

2.3 Trinity Top Ten

The Trinity Top Ten has almost completely collapsed. The level of difficulty in running this big a competition was totally underestimated, and when some clubs did not respond and delayed sending their paperwork, the whole competition was comprimised. The results are still unavailable, and I believe that we cannot run the TTT next year. We should instead concentrate on other aspects of the club. This is regrettable, but it is simply not possible to run a postal competition without having one person whose only job is the TTT, as opposed to having a committee member with other duties assigned to running it.

3 External Competitions

3 External Competitions

3.1 National Level

3.1.1 Air Rifle Nationals 2001

Venue  UCD Rifle Club, Belfield
Date  May 12/13, 2001
DURC Attendees  Mark Dennehy

Ben Jones Keith White

Scores

Mark Dennehy  513
Ben Jones  505, 488

Prizes Won

Ben Jones  3^{rd} Class D

DURC were asked to assist with the logistics for the Nationals by the NTSA on the basis of the standard set by the DURC Air Rifle Open earlier in the season. This was a vindication of the effort we put into the DURC 10m Open last year.

Keith White was the offical scorer and statistician for the competition, as he has been for most of the 10m competitions in the past few seasons. I shot a 513 (ex600) on the first day but due to a foul-up, was disqualified from the Open when my second match had only 55 shots, one of my target cards being lost in the meleé that Sundays shooting proved to be. Ben shot very well, scoring a 505 on his first day, which was the first time he has broken 500. Ben Jones managed to get third place in Class D, beating off virtually all of the Wilkinstown club to do so (some twenty shooters!); and he was promoted to Class C. This was the last air shoot of the 2000-2001 season.

3.1.2 Fassaroe 50m Open 2001

Venue  Fassaroe Shooting Club, Enniskerry
Date  June 16-17, 2001
DURC Attendees  Steven Watterson

Mark Dennehy John Keeney Alan Smyth Joe Lalor Peter Byrne Ray Reilly

Scores

50m Prone Rifle

Ray Reilly  571
Steven Watterson  570
Mark Dennehy  556
Peter Byrne  549
Joe Lalor  534

100yd Prone Rifle

John Keeney  555
Alan Smyth  511
Joe Lalor  338

Prizes Won

John Keeney  First Place, 100yd Prone Rifle

The weather was awful on the Sunday. While we had no entrants for the Three Position match , we had great success in the Scottish Match (100yd Prone Rifle), with John Keeney winning not only the handicapped section but the open section as well, with an excellent score of 555. 50m saw more entrys, but this was Joe and Peter’s first 50m competition, and I was breaking in my new gear, so we took it as a training exercise.

3.1.3 50m Prone and 3P Nationals 2001

Venue  Fassaroe Shooting Club
Date  July 7th, 8th, 2001
DURC Attendees  Mark Dennehy

John Keeney Alan Smyth Adrian Costigan Peter Byrne

Conor McDermottroe

Joe Lalor

Scores

Three Position Rifle

Mark Dennehy  937

Prone Rifle

John Keeney  577
Alan Smyth  569
Ray Reilly  566
Conor McDermottroe  562
Mark Dennehy  561
Peter Byrne  557
Adrian Costigan  554
Joe Lalor  522

Prizes Won

John Keeney  First in Class B, Silver Medal overall
Alan Smyth  First in Class D
Adrian Costigan  Second in Class C
Peter Byrne  Third in Class D
Joe Lalor  Third in Class C

DURC entered a reasonable contingent in this year’s 50m Prone Nationals and all did well. John Keeney managed to score 577 and take the overall second place, making up for Ray Reilly losing the title of National Champion somewhat. Alan Smyth also shot well, making it to the 10-shot final and placing 7th to do so. Newcomers to the 50m sport did well also, with Adrian Costigan, Joe Lalor, and Peter Byrne walking away with prizes, though Conor McDermottroe was only denied a prize by a contraversial classification decision.

Only one competitor entered the three-position match however, and I did not shoot as well as I had hoped, coming in fourth but increasing my score by about a hundred points from my last 3P match. DURC was planning on training for 3P in earnest this year, using Air Rifles due to range restictions, but time pressures on the committee mooted this idea.

3.1.4 DURC 50m Open 2001

Venue  Fassaroe Sporting Club
Date  July 22
DURC Attendees  Steven Watterson

Ray Reilly Peter Byrne Alan Smyth Mark Dennehy John Keeney Adrian Costigan

Scores

Steven Watterson  585
Ray Reilly  582
Peter Byrne  577
Alan Smyth  572
Mark Dennehy  568
John Keeney  566
Adrian Costigan  545

Prizes Won

Steven Watterson 1st place overall, 2nd place class A, 2nd place Rapid-Fire
Ray Reilly  2nd place overall,3rd place class A, 1st place rapid-fire
Adrian Costigan  2nd place class C
Alan Smyth  1st place class D
Peter Byrne  2nd place class D

Steven Watterson won the Dublin University Rifle Club Open, beating Olympian Alan Lewis with a score of 585. Steven had just returned from the Island Games in the Isle of Man where he won two silver medals in Air Rifle and 50m three-position Rifle, as well as a bronze medal. Steven won a custom rifle barrel kindly donated by Shilen ltd. Second place went to Ray Reilly and third to Tony Cogan who travelled to Zagreb later that week to represent Ireland in the European Championships.

The Rapid-Fire competition was won by Raymond Reilly, with an impressive 97/100. Second place was decided by a shoot-off between Steven Watterson and Alan Lewis, and was won by Steven Watterson.

3.1.5 Comber Open 2001

Venue  Comber Rifle Club, Comber
Date  August 4-5, 2001

DURC Attendees

Steven Watterson  Mark Dennehy
John Keeney  Peter Byrne

Ray Reilly

Scores

100 Yards

Steven Watterson  384
Ray Reilly  380
John Keeney  383
Peter Byrne  384
Mark Dennehy  342

50 Metres

Steven Watterson  577
Ray Reilly  562
John Keeney  570
Peter Byrne  569
Mark Dennehy  536

Teams Competition

”DURC Angles”

Mark Dennehy  567
John Keeney  583
Peter Byrne  578
Ray Reilly  579
Total  2307

Third Place

”North&South”

Steven Watterson  581
Total  2332

First Place

Prizes Won

Steven Watterson  First place, teams competition

Eight place overall

John Keeney  3rd, Class B, 50m
Peter Byrne  2nd, Class B, 100yds

The Comber Rifle Club is renowned throughout Ireland and the UK as being the toughest range in either island to shoot on. Situated so close to the shore, tidal winds are an enormous challange, and the topography of the range means that wind conditions can change from moment to moment and one firing point can have a totally different wind direction from another only a few feet away. As a result, the 50m Open held in Comber each year is perhaps the most challenging Ireland has to offer and it is little surprise that shooters from Comber such as Olympian Gary Duff, are heavily featured amongst Ireland’s best.

Comber Rifle Club

DURC competeted in this year’s Comber Open with a squad of five shooters. Logistics proved to be a stumbling block, with far too much equipment for one car. The conditions were typical for Comber - changing winds, changing light levels and occasional torrents of rain.

All of the DURC shooters acquitted themselves well, except perhaps for myself, though I had more than my fair share of problems - the sights for my rifle were accidentally not packed so I had to jury-rig a sights setup, and to add insult to injury I was bitten by some insect and had an allergic reaction to the bite.

Steven managed to display a total lack of impediment however, and made it to the finals round, though he didn’t place for a medal.

John Keeney and Ray Reilly

Steven Watterson reflecting between cards during the finals

The finals; and John and Steven relaxing after a long day of shooting

3.1.6 East Antrim Festival 2001

Venue  East Antrim Rifle Club
Date  7-9 October, 2001
DURC Attendees  John Keeney

Ray Reilly Alan Smyth

Scores

10m Air Rifle

John Keeney  477
Alan Smyth  485

10m Air Pistol

John Keeney  426
Alan Smyth  477

50m Prone Rifle

John Keeney  543
Ray Reilly  575
Alan Smyth  540

25m Smallbore Pistol

John Keeney  303
Alan Smyth  427

Following the successful Freshers’ Week membership drive, three members of DURC attended the EARC Festival. Due to complications with paperwork, a change of dates and the level of work during Freshers’ Week, our original planned team of eight was seriously whittled down.

John and Alan shot very disappointing scores due to extreme lighting conditions early saturday morning, where the early morning sun was directly behind the target thereby eliminating almost all contrast in the sight picture.

It’s hoped that next year we will be able to send more people to EARC for the competition, this year saw a lot of factors come together to prevent this.

3.1.7 Wilkinstown Open 1 2001

Venue  Wilkinstown Target Shooting Club
Date  14 October 2001
DURC Attendees  Steven Watterson

Mark Dennehy Ben Jones Peter Byrne

Scores

Steven Watterson  571
Mark Dennehy  544
Ben Jones  485
Peter Byrne  453

Prizes Won

Steven Watterson  3rd Place, Class A

3rd Place Overall

Mark Dennehy  2nd Place, Class B
Peter Byrne  1st Place, Class C

DURC sent a team of four to the first competition of the 2001-2 Air Rifle season. Attendance was low due to difficulties with transport and lack of training time for new club members.

The new WTSC mobile range was used: this range is currently the largest in the country with 15 firing points. It currently hosts the National Squad training sessions as well as other large competitions. Previously, only the UCD range could cope with these events.

The DURC contingent arrived in WTSC early on Sunday morning, despite having attended the Olympic Council of Ireland’s seminar on coaching in the University of Limerick the day before. Mark, Peter and Ben shot on the first detail, Steven acting as “buddy” for all three. The “buddy” system, recently introduced in DURC by Steven, has proven a great success - the “buddy” acts as another pair of hands for the shooter while setting up on his/her firing point and general dogsbody during the match to minimize disruption for the shooter, who then trades role with the “buddy” on the next detail. This allows for one-to-one supervision of new shooters as well as improved performance of experienced shooters.

Steven and Mark both set personal bests of 571 and 544 respectively and both went through to the ten-shot final, where Steven had the best final score of the match. Mark was shooting with his new air rifle and other gear - and despite some silly technical errors that lost him up to 15 points, shot a PB. Proving that the right equipment can make a serious difference to performance. For reference, 570 is the Minimum Qualifying Score, or MQS, for the Olympics; 537 is the qualifying score for the National Developmental Squad; and 559 is the qualifying score for the National Squad.

3.1.8 UCD Open 1 2001

Venue  UCD Rifle Club, Belfield
Date  28-29 October 2001
DURC Attendees  Steven Watterson

Mark Dennehy Ben Jones Peter Byrne Adrian Costigan Nithin Santhosh Dave Barron

Colman Marcus-Quinn Ljiljana Skrba Daniel O’Neill

Scores

Steven Watterson  571 (669.8)
Mark Dennehy  550
Ben Jones  491
Peter Byrne  475
Adrian Costigan  474
Nithin Santosh  430
Dave Banon  377
Colman Marcus-Quinn  359
Ljiljana Skrba  321
Daniel O’Neill  281

Prizes Won

Steven Watterson  Third Place Overall

Third Place, Class A

Mark Dennehy  Third Place, Class B
Peter Byrne  Second Place, Class D
Adrian Costigan  Third Place, Class D

This was the first Air Rifle competition of the season that we can get large numbers of people to, owing to the nearby location of UCD. We had ten people attend in total, breaking down to five experienced shooters and five novices, a reasonable demographic.

Steven, Mark, Peter and Adrian all shot personal bests, and secured prizes. Mark, at 550, is now set to enter class A after the next competition. Nithin, Dave, Colman and Ljiljana all shot very well for their first competition, especially considering that they had only had two or three half-hour details of shooting prior to the competition.

3.1.9 Leinster Indoor Open

Venue  Rathdrum Rifle and Pistol Club
Date  November 4th, 2001
DURC Attendees  Mark Dennehy

John Keeney Peter Byrne Deirdre Ninaber Adrian Costigan Eric Farrell

Michael Madigan Shane Kelly Alan Smyth

Scores

John Keeney  581
Alan Smyth  578
Deirdre Ninaber  568
Peter Byrne  564
Adrian Costigan  550
Eric Farrell  550
Mark Dennehy  549
Shane Kelly  535
Michael Madigan  518

Prizes Won

John Keeney  Second place overall

First place, Class B

Alan Smyth  First place, Class C
Deirdre Ninaber  First place, Class D

DURC sent a squad to Rathdrum for the Leinster Indoor Open. The early start took it’s toll on some - I shot the worst score I have shoot indoors since I began competing eight years ago. John Keeney, however, proved far more resiliant and despite having travelled from Kildare only the night before, shot a 581 to take second place overall, beaten only by Dave Cullen, a 40-year veteran of target shooting, by a mere 5 points and on Dave’s home range too! John also took the first place prize in class B.

In fact, DURC swept the boards in the classification prizes - first place in Classes B, C and D went to DURC shooters John Keeney, Alan Smyth and Deirdre Ninaber. Deirdre was competing in her second competition ever - having only ever shot before in the 25yd Nationals, scoring a 491. Therefore in one competition she has improved by 77 points - an incredible leap in performance which hasn’t been seen in DURC in many years.

John Keeney and Deirdre Ninaber accepting their prizes

3.1.10 Fassaroe 10m Open 1 2001

Venue  UCD Rifle Club
Date  10-11 November 2001
DURC Attendees  Steven Watterson

Mark Dennehy Ben Jones John Keeney Peter Byrne Adrian Costigan

Coleman Marcus-Quinn

Scores

Steven Watterson  565
Mark Dennehy  553, 552
Ben Jones  486
John Keeney  438
Peter Byrne  476, 465
Adrian Costigan  457, 427
Coleman Marcus-Quinn  441

Prizes Won

Mark Dennehy  Class B, 1st place
Peter Byrne  Class D, 3rd place
Coleman Marcus-Quinn Best DURC Novice shooter

DURC sent a small team to compete in the Fassaroe Club’s Air Rifle Open (aka the WAIDAM open), with a total of 7 competitors.

The standard of competition was high, the winning score being 578 from Nicola Murphy of Wilkinstown (an ex-UCD shooter), who recently won silver in the English championship. Steven shot lower than his previous average, with his 565 beaten into 4th place by one point by Declan Kelly of the Fassaroe Club (566). (Second place was taken by Rhona Barry with 570).

Unfortunately, most of the DURC shooters were only just recovering from their annual bout of ’flu and so scores were generally down. I shot a personal best on Saturday with 553 and then reentered on Sunday hoping to push that further, but instead dropped one point to 552. Peter Byrne shot lower than he had hoped, but still managed to take third place in class D. Novice Colman Marcus-Quinn managed to increase his Personal Best score by 82 points to 441 and took the prize for best DURC novice.

3.1.11 DURC 10m Open 2001

Venue  UCD Rifle Club, Belfield
Date  24-25 November 2001
DURC Attendees  Steven Watterson

Mark Dennehy Ray Reilly Liam Murray Alan Smyth Peter Byrne Adrian Costigan Dave Barron

Victoria McDowell

Scores

Steven Watterson  565
Mark Dennehy  554,548,533
Ray Reilly  535
Liam Murray  502
Alan Smyth  497
Peter Byrne  467,500
Adrian Costigan  456
Dave Barron  382
Victoria McDowell  363

Prizes Won

Steven Watterson  Third place overall
Liam Murray  First place class D
Peter Byrne  Third place, Juniors

The annual DURC 10m Air Rifle Open is one of our two most important external competitions. Attendance was high, with 35 entrants and some 43 entrys in total, from clubs as far away as Antrim. The shoot was a registered event, which means it is one of the 6 air rifle events during the year where the scores may qualify a shooter for selection to represent Ireland internationally.

The standard of competition was higher than it has been in many years, with all eight finalists scoring above 550 in the qualifying round. Two DURC members, Mark Dennehy and Steven Watterson made it to the finals. Mark placed in last place in the finals. Steven went into the finals tied for third place with Junior Squad shooter Richard Stapleton on 565 points. A very tight finals ensued, with Richard and Steven exchanging places with nearly every shot. Steven managed to fend off Richard with his last two shots however, to take third place by 0.3 points on 662.9 points (out of a possible 709.0).

John Keeney acting as Range Officer during the ten-shot Finals, with Ben Jones acting as Statistics Officer

The finals saw three such ties decided - Declan Kelly and Robert Clarke started on 561 and exchanged places several times, with Declan finally winning out by 0.7 points. Rhona Barry and Nicola Murphy both started on 574, tied for first place, but Nicola proved to be the stronger shooter on the day and won by 1.1 points in total with a final score of 673.5. Nicola was in fact the story of the day - a trigger malfunction had caused her to lose 9 points in one shot during the qualifying round - that she was able to put this disaster firmly behind her and go on to tie for first place was an inspiring performance.

Other DURC members did well also, with Liam Murray breaking 500 for the first time and claiming the first place in Class D, even though he couldn’t attend the award-giving ceremony. Peter Byrne also set a personal best, hitting 500 right on the nose. Alan Smyth just missed the 500 mark by 3 points. Ben Jones missed it by 6 with 494. Adrian Costigan put in scores of 467 and 456 repectively. Newcomers Dave Barron, Steven Beck and Victoria McDowell also did admirably, though they did not reach the rostrum. And Ray Reilly made a welcome return to the Air Rifle scene after a prolonged absence caused by a back injury. He put in a respectable 535.

Keith White, though he did not shoot, proved invaluable yet again as a Statistics Officer, responsible for scoring targets and tracking peoples scores during the competition, which he did using his own custom software which is due for general release Real Soon Now.

Mr. Terry McAuley, the DURC president and Director of Sport for Trinity College, graciously agreed to come to the Open and present the prizes. Fortunately, he was able to observe the 10-shot Finals and was suitably impressed by the high level of shooting. He gave a short speech on the plans for the new Sports Center and it’s planned Rifle Range, and then presented the prizes. We’d like to thank Mr.McAuley for donating his time to the club for this event, as it certainly made an impression on the target shooting community.

We would also like to thank Hammerli, who sponsored us with some baseball caps for all the Junior shooters at the event, and especially Mr. Simon Kennedy, an ex-Captain of DURC, who gave a large contribution to the prize fund. As Simon was a member of the National Air Rifle Squad for many years and represented Ireland internationally in Air Rifle, we felt that the commissioning of the Simon Kennedy Air Rifle Shield was appropriate: it will be awarded each year to the overall winner of the DURC 10m Air Rifle Open.

Director of Sport Terry McAuley presents the Simon Kennedy Air Rifle Shield to the overall winner Nicola Murphy, pictured here with Runner-up Rhona Barry and bronze medallist Steven Watterson.

3.1.12 UCD Open 2 2002

Venue  UCD Rifle Club, Belfield
Date  23-24 February 2002
DURC Attendees  Steven Watterson

Mark Dennehy Ben Jones

Nithin Santhosh Rob McAdoo Ljiljana Skrba Kate Fanning

Victoria McDowell

Scores

Mark Dennehy  541
Ben Jones  504
Nithin Santhosh  504
Kate Fanning  442
Victoria McDowell  413
Ljiljana Skrba  382

Prizes Won

Ben Jones  Second Place, class C

Our 40th anniversary celebration took place the night before, and unfortunately all this celebrating meant that our numbers were depleted, as many people were still asleep when thy should be shooting!

Nithin Santhosh shot a personal best of 504 and earned promotion to class C, breaking the 500 barrier for the first time. Ben Jones equalled Nithin’s score (after two successful challanges when shots were mis-scored). I reached the final and moved up to sixth place when John Woodward of WTSC had a disasterous first shot. Steven Watterson was unable to shoot as supervisory duty claimed the first day of competition and the DURC anniversary claimed the second - a shame, as given Steven’s recent scores and those put in on the weekend, we probably missed out on second place in the overall competition.

Kate Fanning, Victoria McDowell and Ljiljana Skrba all shot well and improved on their personal best scores. All this means we’re going to have an excellent team for the colours match against UCD, which is taking place on April 27th.

3.1.13 Wilkinstown Open 2 2002

Venue  Wilkinstown Target Shooting Club, Wilkinstown
Date  March 9-10, 2002
DURC Attendees  Mark Dennehy

Kate Fanning Ray Reilly Keith White

Scores

Mark Dennehy  555
Ray Reilly  547
Kate Fanning  383

Despite inclement weather and horrible driving conditions, DURC sent a small team to the WTSC Open. Shooting conditions were less than optimal as the weather caused the light in the range to vary from bright to dull, which of course upset the sight picture for the shooters. I managed to compensate, and set a new personal best of 555 points, and placed 6th in the final, fighting off Helen Meade for the second competition in a row. Ray Reilly also showed progress back to his original form with a 541.

And finally, Keith White’s software for administering competitions is nearing it’s first release. Allready it has become the standard software for this purpose in Ireland, and he should be congradulated on this as he has put nearly 8 years of development into it, with extensive user testing.

3.1.14 Air Rifle Nationals 2002

Venue  Wilkinstown Target Shooting Club, Navan, Co.Meath
Date  April 13/14, 2002
DURC Attendees  Steven Watterson

Mark Dennehy Ray Reilly Keith White Peter Byrne Nithin Santhosh Adrian Costigan Kate Fanning John Keeney Rob McAdoo

Victoria McDowell Doug Woolett Deirdre Ninaber

Ben Jones

Scores

Steven Watterson  574,578
Mark Dennehy  543,538
Ray Reilly  544,519
Peter Byrne  531,528
Nithin Santhosh  526,524
Adrian Costigan  488,522
Kate Fanning  491,500
John Keeney  506,483
Rob McAdoo  457,422
Victoria McDowell  426,419
Doug Woolett  378,427
Deirdre Ninaber  537
Ben Jones  533

Prizes Won

Steven Watterson  Bronze, National Open; 3rd, Class A
Peter Byrne  2nd Class C, National Open; 3rd Class C, National Championships
Nithin Santhosh  3rd ClassC, National Open
Kate Fanning  1st Class D, National Open
John Keeney  2nd Class D, National Open; 3rd Class D, National Championships
Rob McAdoo  3rd Class D, National Open
Deirdre Ninaber  1st Class D, National Championships
Ben Jones  2nd Class C, National Championship

DURC sent a large team of 14 people (13 competitors and one administrative officer) to this year’s Air Rifle National Championships, which were held in Wilkinstown this year for the first time in a decade, in the new NTSA portable range. This range shows how uncomplicated the requirements are for an olympic discipline like 10m Air Rifle. The full range seen here can be disassembled and moved in only three hours, though it does require several people and vans to move the various components. So long as an appropriately sized venue is available, the range can be used. This range should be examined with the new projectile design

in mind - it demonstrates how the most important requirement is the physical size of the hall, but also the necessity for permanent fixtures - a three hour setup time would be prohibitive in anything but a temporary range like this one.

The DURC squad shot extremely well, even given the required 5am starts on Saturday and Sunday (necessary to arrive in Wilkinstown with sufficent time to prepare to shoot). All bar myself set personal best scores, and Steven held the lead in the National Open until the last few shots. The National Open was held concurrently with the National Championship to allow for international shooters like Steven to compete. Steven managed to secure the bronze medal for DURC in the Open. He would have taken a national championship medal home, but for his passport.

The Colours squad used this event as a warm-up and training exercise for the colours match (due to be held in Wilkinstown on April 27th), and given the scores acheived, this was judged to be a resounding success. The squad learnt many valuable lessons in match strategy, and also gained some experience of shooting under pressure. Deirdre Ninaber, shooting in her first air rifle match, managed to take first place in class D with a 537, a phenonomally high score for a beginner, and which will see her moved up to class B with her next competition. This bodes very

well for the upcoming colours match, which this year is an offically CUSAI–recognised Intervarsity for the first time.

The DURC Squad. Back Row, left to right: Doug Woolett, Ray Reilly, Peter Byrne, John Keeney, Adrian Costigan, Ben Jones, Mark Dennehy Front Row, left to right: Victoria McDowell, Deirdre Ninaber, Kate Fanning and captain Steven Watterson

3.1.15 Colours

Colours is no longer simply the UCD/TCD colours match of old. CUSAI finally offically recognised the competition this year, and so this year is the first Irish 10m Air Rifle Intervarsity. The colours match remains, as a seperate sub-competition between us and UCD, at least for the moment. Final teams have not yet been selected, but this year has seen the most serious and organised training to date for this event, and I’m happy to predict a thorough stomping for UCD, with us taking the Senior and Junior shields back.

3.2 International

3.2.1 Island Games 2001

The first Island Games were held in 1985 as part of ’The Year of Sport’ in the Isle of Man. The intention was to bring together sportsmen and sportswomen from a number of small islands from different parts of the world to participate in a festival of sport. So from 14 islands - Aland, Anglesey, the Faroes, Froya, Gotland, Guernsey, Hitra, Iceland, Jersey, Malta, the Orkney Islands, the Shetland Islands, St. Helena and the Isle of Wight - athletes came to the Isle of Man to compete in 7 sports - badminton, athletics, cycling, shooting, swimming, soccer and volleyball. The ’festival’ was such a success that at its conclusion the Island Games Association (IGA) was instituted and the decision was taken to hold Island Games every two years in different member islands.

This year Steven, who is a Manx native, was competing, and did spendidly well. In 10m Air Rifle, he placed third in the Individual event, shooting his first offical ’tun’. We’re still not sure if this is a club record though, because he wasn’t shooting to represent DURC at the time. Despite a setback of shooting a 7.0 on his first shot in the final, he fought hard and managed to fend off David Turner and Pierre Goubert for third place. In the Team event, Steven and teammate Jaccie Hamilton-Jones came fourth, beaten by the Gotland, Jersey and Aland teams - though Jaccie had some problems with her air rifle causing her to shoot a very low score of 493 to Steven’s 554.

In 50m 3P rifle (the most rigorous and demanding of all rifle disciplines), Steven managed to take the silver medal, beaten only by Mattias Magnusson of Gotland. It was a close race for the qualification round, as Stevens 1075 beat Birger Johansson of Gotland into third place by only one point. However, in the final Steven widened the gap to 12.4 points, taking second place with no problems whatsoever. In fact, his 89.7 was the highest score shot in the final round. In the team match, Steven and teammate Barry Jones took second place, beating the Jersey team off by only eleven points at 2064 to 2053, thus securing Steven’s third medal of the Games.

3.2.2 Munich World Cup

The Munich World Cup was attended this year by 65 nations with nearly 650 shooters competing. Ireland sent a team of two this year, Rhona Barry (who competed in the Atlanta Olympic games) and Alan Lewis (who competed in the Sydney Olympic games). Mark Dennehy, was sent to act as Team Manager.

Alan’s competition, Mens 50m Prone Rifle, did not go well for him and he missed the cutoff in the elimination round by one place, with a score of 583 (584 being high enough to qualify for the next round).

Rhona did not have an elimination round in her event, women’s 10m Air Rifle. She shot a 383 in qualification, which is a Minimum Qualifying Score (MQS) for the Athens Olympic Games. This makes her the first Irish shooter to qualify for selection to fill a quota place for the games, and also one of the first Irish athletes to do so.

Mark was present in an administrative capacity during the event, and also represented the Irish contingent at the Technical Meeting and the Delegates Meeting during the World Cup.

3.2.3 Isle of Man Easter Shoot

The Isle of Man Easter shoot was supposed to be a major event this year. Last year’s “fact-finding” mission provided the data on organisation and logisitics, UCD and WTSC were all sending teams and it was going to be a large event. Sadly, however, interest fell apart in the last few days, with noone willing or able to spend the cash on attending, and with the GB Juniors training taking members of both WTSC and DURC’s teams. Next year, we should be promoting this a bit more as it is an idea small-scale multi-disciplinary international event.

3.2.4 Bisley Week

John, Ben, Alan and Ray attended Bisley Week this year, and happily promised a write-up. In August 2001. Which I still don’t have, though we do have the photos, which are on the website. And we know Ben took fourth in the NSRA/Eley Postal Finals, which was impressive.

3.3 Postal Competitions

Postal competitions represent the easiest and cheapest way for DURC to compete internationally, and as such are vital to the club’s competitive edge. This year we have taken on summer season postals along with the traditional winter UCESSA ones, to keep the 10m people sharp during the “off-season”.

3.3.1 UCESSA 25yd Prone Rifle

This competition saw it’s usual share of late shooters, cards filled out incompletely and missed by me before posting. However, we had an exceptionally good year aside from this, with almost everyone who shot all their cards coming in the top three of their class. First places were taken by John Keeney, Alan Smyth and Deirdre Ninaber, seconds by Peter Byrne, myself, Adrian Costigan and Michael Madigan. Special note to Deirdre, who won nine out of ten rounds for a nearly-perfect points score of 99/100.

Division Name  Score  Place
1  2  3  4  5  6  7  8  9  10 Total
2  Reilly R.  NCR 99 98 99 98 95 97 96 98 99  879 7th
2  Watterson S.  99 99 97 dq 97 98 97 97 100 100 884 6th
3  Keeney J.  96 99 99 100 99 98 95 98 98 99  981 1st
5  Byrne P.  99 97 95 100 97 96 97 98 99 98  976 2nd
7  Dennehy M.  97 99 98 98 95 95 97 96 96 95  966 2nd
9  McDermottroe C. dq 97 97 96 98 97 94 99 NCR NCR 678 5th
14  Smyth A.  dq 97 97 98 98 96 97 96 95 99  873 1st
16  Costigan A.  95 92 98 89 96 95 94 97 96 96  948 2nd
19  Murray L.  94 92 NCR 91 87 97 94 NCR NCR NCR 555 7th
21  Ninaber D.  99 97 96 95 94 98 91 94 96 94  954 1st
21  Madigan M.  NCR 94 95 88 95 95 91 93 96 91  838 2nd
21  Gallagher D.  NCR 86 NCR NCR NCR NCR NCR NCR NCR NCR 86 8th
Division Name  Points  Place
1 2 3 4 5 6 7 8 9 10 Total
2  Reilly R.  0 8 6 9 6 4 5 3 5 9  55 7th
2  Watterson S.  7 8 3 0 4 6 5 4 10 10  57 6th
3  Keeney J.  6 10 10 10 9 9 4 8 9 10  85 1st
5  Byrne P.  9 8 4 10 10 4 6 9 10 9  79 2nd
7  Dennehy M.  10 10 10 10 4 6 8 10 9 7  84 2nd
9  McDermottroe C. 0 9 9 7 10 9 3 10 0 0  57 5th
14  Smyth A.  0 10 10 10 10 10 10 10 8 10  88 1st
16  Costigan A.  10 7 10 1 10 9 8 10 10 10  85 2nd
19  Murray L.  9 10 0 6 5 10 10 0 0 0  50 7th
21  Ninaber D.  10 10 10 10 9 10 10 10 10 10  99 1st
21  Madigan M.  0 9 9 8 10 9 10 8 10 9  82 2nd
21  Gallagher D.  0 6 0 0 0 0 0 0 0 0  6 8th
Division Name  Rank
1 2 3 4 5 6 7 8 9 10 & Final
2  Reilly R.  8 9 7 6 6 6 7 9  7
2  Watterson S.  4 2 6 9 9 8 8 9 7  6
3  Keeney J.  5 2 1 1 1 1 1 1 1  1
5  Byrne P.  2 2 4 2 2 3 4 4 3  2
7  Dennehy M.  1 1 1 1 2 2 2 2 2  2
9  McDermottroe C.  8 6 4 4 4 4 4 4  5
14  Smyth A.  1 5 3 2 2 1 1 1  1
16  Costigan A.  1 7 1 4 2 2 2 2 2  2
19  Murray L.  2 1 5 4 5 3 2 5 7  7
21  Ninaber D.  1 1 1 1 1 1 1 1 1  1
21  Madigan M.  7 5 4 3 3 2 2 2  2
21  Gallagher D.  8 8 8 8 8 8 8 8  8

3.3.2 UCESSA 10m Air Rifle This hasn’t finished yet, but Steven’s tearing off ahead as usual, currently holding in fourth place with an aggregate score of 29 points and an aggregate gunscore of 759.

4 Internal Competitions

3.3.3 NSRA 10m Team

We have entered this competition to keep the air rifle section of the club active. Shooting has not yet started, though entries are closed. DURC competitors in this competition are :

Team 1 : Mark Dennehy, Steven Watterson and Ben Jones.

Team 2 : Nithin Santhosh, Ljiljana Skrba and Rob McAdoo.

Team 3 : Ray Reilly, Kate Fanning and Peter Byrne.

Good luck to all.

3.3.4 NSRA 10m Individual

We have entered this competition to keep the air rifle section of the club active. Shooting has not yet started, though entries are closed. DURC competitors in this competition are :

Steven Watterson Mark Dennehy Ray Reilly
Ben Jones Peter Byrne Ljiljana Skrba

Nithin Santhosh

Good luck to all.

4 Internal Competitions

4.1 Ladders

The ladders have seen a fair amount of activity over the past year. The closing date for the ladders was set at Friday, April 19. This year saw the return of the advanced classes in 10m and 25yd, for the first time in some years. Hopefully they’re here to stay!

4.1.1 25yd Prone Rifle

In the advanced class, Steven Watterson took first place with 98.714 points as an average, setting a new club record while doing so. Second place on 97.769 went to John Keeney, beating out Ray Reilly’s 97.250 which took third. Some results have been omitted from this ladder, however, including Ray’s NSRA/Eley tun which also made it to the record books this year. There were 74 cards on the ladder from 5 shooters, not an unreasonable activity level.

In the experienced class, Peter Byrne’s 97.1 took first, followed by Conor on 96.2 and Deirdre on 95.071. 174 cards made it onto the ladder in this section, showing the level of activity.

In the novice class, Brian Long’s 91 beat out Ghandi Solayar’s 90.9 by a tight margin to take first place, with W.Sim’s 87.125 taking third. Activity was not so high here, with only 30 cards on the ladder in total.

4.1.2 10m Air Rifle

119 cards from three shooters made up the advanced ladder, an startlingly high level of shooting. First place was, without surprise, Steven Watterson, with 95.2 as an average (some of us count that as our personal best, dammit). I got second place with 92.4 and Ray is making a good comeback to 10m with 91.1 as his average.

In the experienced class, Peter and Ben had a scrap to see who could shoot the most cards, Peter winning with a staggering 97 cards shot. (that actually 194 5-strip targets...) Deirdre, however, took first place with 87.893 to Peter’s second-place 87.494, with Ben taking third place on 86.4. All very close indeed, and with a total of 330 cards on the ladder, we can be absolutely sure that the results are no fluke.

Novice class wasn’t terribly quiet by comparison, 269 cards from 15 people made it the most crowded, if not the busiest of the ladders, but the level of training belies the shooters’ novice status, as most of the colours squad is hiding in here doing dry-firing to train. Nithin the incorrigible took first place with 82.889, with Ye “Elena” Wang taking second on 78.636 and Kate Fanning in third place on 76.8. Nithin also shot the most cards, with 43 compared to the next nearest total of 35 (shot by both Kate and Victoria McDowell).

4.1.3 25yd 3P Sporter Rifle

Sporter was very much the “try it once and walk away” discipline this year - of the 11 people on the ladders, only 3 people had the necessary three cards to qualify for a prize. In fact the top two on the ladder, myself and Steven (125 each), didn’t shoot enough cards and so the first place goes to Adrian Costigan, with 123.333. Second place skips John Keeney’s single 123 entry to go to Frank De Salle, who shot the most cards (8) and got an average of 122.429. Third place goes to Mick Durcan, who shot an average of 117 and 6 cards in total, skipping Ben and Michael Madigan. A total of 29 cards were shot, an extremely quite discipline. We need to get some people going to external competitions with these rifles, wheither in Silhouette or Sporter competitions, or we risk the discipline dying out altogether.

4.2 Championships

The championships this year were not as well-organised as last year, though they did succeed better on the social gathering side of things. Next year’s committee are strongly urged to set the calendar for events before the start of michaelmas term, and to announce it widely well in advance.

4.2.1 25yd Prone Rifle

25yd proved contraversial this year, with three people tied on 98 for first place. After much arguing as to whether a pip had to be clipped or obliterated to count, graduated guages were instead used to rescore the cards. Following this, Steven took first place, Adrian second and Ray took third place. Following this, it should be suggested that next year the rules be written down *before* the event, and that anyone who thinks they may have first place shouldn’t join the party in case a shoot-off is called for!

4.2.2 10m Air Rifle

Air Rifle proved interesting this year, with equipment limitations for the first time (no trousers for anyone). Kate Fanning took first place, trouncing all comers with an 86, while Peter, Frank and Adrian all tied on second place with 80. Counting tens to break the tie, Peter took second with 3 tens, Frank takes 3rd with 2, and Adrian 4th with 1.

4.2.3 3P Sporter Rifle

3P was well attended, being the first internal competition and while John Keeney left early with a 123, happily convinced of a win, Steven and Mark stomped all over this with Steven taking second with 125 and Mark taking first with 130.

4.2.4 Rapid Fire

Rapid Fire was again interesting, with myself getting a really disappointing 85 (with a -10 on the last shot) - which would have been less crushing if my scope hadn’t been pointed at Steven’s card...

First place went to Alan with 93, second to Conor on 92 and Steven took third on 91.

5 Club Records

4.3 Other

4.3.1 Standards

Disappointingly, noone shot for standards this year, despite the interest shown in the NSRA version of the standards. I would urge anyone who is shooting regularly to try for the standards, and anyone on the colours team should definitely try for them. Anyone interested should contact me or the next Match Secretary.

4.3.2 GB Team stuff

Though they haven’t happened yet, and they’re not open to DURC shooters, we ought to note that Steven’s going to the Commonwealth Games on the GB team. These are at a level somewhere between the World Cups and the Olympics, and it’s quite an achievement, especially given the high standards of the GB team.

And while we’re at it, also note that Peter Byrne was invited to go training with the GB Juniors this easter, which is the first time we’ve had someone do that in many years.

5 Club Records

This year sees a reorganisation of the Club records by discipline. It also sees the introduction of the 3P 25yd Sporter Records, as we have not shot 4P in some time. Next year should see a more standardised approach to club records, with individual divisions having standard records - Best Individual Card, Highest Ladder Average and Best Competition Score and Best Competition Placing for all. Also, record equalling should be tracked.

5.1 25yd Prone Rifle

Best Individual Card

Andy Trayler 1993 100, 9 pips shot out
Ray Reilly  2002 100, 9 pips shot out

Highest Ladder Average

Ray Reilly  1999 98.231
Steven Watterson 2002 98.714

This displaces Ray Reilly’s previous record of 98.231, set in 1999. However, some records arer missing from the ladder, so this has yet to be verified.

Highest Competition Score

Score Ray Reilly 1999 595

(Rathdrum Autumn Open)

Placing Mark Dennehy  2001 Gold Medal, National Championships

Highest Rapid-Fire Score

Derval Tubridy 1993 98

5 Club Records

5.2 100yds Prone Rifle

Best Individual Card

100yds William Walsh 1987 200(ex200)

5.3 50m Prone Rifle

There is no Best Individual Card division, due to a lack of records. My successor should endeavour to locate records to establish who holds this record. Highest Competition Score

50m Prone Ray Reilly 1999  593

(EARC Festival)

Placing Ray Reilly 2000  Gold Medal, National Championships

5.4 50m 3P Rifle

Highest Competition Score

50m 3P mens Steven Watterson 2002 Easter Shoot IoM 1093
50m 3P ladies Derval Tubridy  519

This displaces Simon Kennedy’s 1995 record of 1084 for men’s 3P.

5.5 25yd 4P Sporter Rifle

Best Individual Card

4P, Sling Keith White  1993 114(ex120)
4P, No Sling Simon Kennedy 1993 107(ex120)

Highest Ladder Average

Conor Tighe 1994 107

5.6 50m 4P Sporter Rifle

Highest Competition Score

Conor Tighe 1994 160(ex200)

5.7 25yd 3P Sporter Rifle

This is a partially new section, as the 3P records started in 2001. Highest Ladder Average

Frank De Salle 2001 126.25

Highest Competition Score

Kenneth Suzuki 2001 134(ex150)

5.8 Prone Sporter Rifle

We have not shot this in many years apparently. Highest Competition Score

25yd Simon Kennedy 1994 384(ex400)
100m Alan Murphy  1990 363(ex400)

5.9 10m Air Rifle

Best Ten-Shot Series

Conor Tighe  1995 99
Steven Watterson 2002 99
Steven Watterson 2002 100

Steven’s tun was shot while representing the Isle of Man in the Island Games. As such, we’re not sure if it should be a DURC record, but a tun in air rifle is worth remembering. Highest Ladder Average

Steven Watterson 2002 95.2

This displaces Simon Kennedy’s 1997 record of 94.467. Highest Competition Score

60-shot Steven Watterson 2002, National Open 578
40-shot Derval Tubridy 1997 380

This displaces Ray Reilly’s previous 575 record from 1999.

5.10 Fullbore Rifle

Highest Competition Score

200yds M.G.Notley  1963 48(ex50)
300yds M.C.C.Heaton 1962 49(ex50)
500yds P.G.Coulson  ?  48(ex50)
P.Gardiner  ?  48(ex50)
M.G.Notley  ?  48(ex50)
600yds William Walsh ?  48(ex50)
900yds C.Day  ?  45(ex50)
1000yds C.Oakley  ?  45(ex50)

Note that fullbore records date circa 1960 and are spotty at best!

5 Club Records

5.11 Miscellaneous Records

3P Air Rifle Andy Traylor 1992  278(ex300)
Triathlon Conor Tighe 1994  298(ex300)
University Top XX Albert Jordan 1998  7 points lost

first Irish shooter to win Individual Gold Medal

My successor should tidy up this section.